TextConv

An Amiga,Unix,Macintosh,MS-DOS,Windows,OS/2,Atari ASCIItext Converter

English version 1.51 Mailware \begin{matrix}c \\ ◯ \end{matrix}1995-97 SciTech Software Development

CONTENTS

Contents

1 General 3

2 Introduction 4

3 System requirements, compatibility 6 3.1 Configurations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 Additional notices . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Warranty, mailware, copyright, copying 8 4.1 Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2 MailWare . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.3 Copyright, copying . . . . . . . . . . . . . . . . . . . . . . . . . . 8

5 Installation 10 5.1 Standard installation process . . . . . . . . . . . . . . . . . . . . . 10 5.2 Additional installation notes . . . . . . . . . . . . . . . . . . . . . 10

6 Starting of TextConv 11

7 Using TextConv 12

7.1 Description of the gadgets . . . . . . . . . . . . . . . . . . . . . . 13 7.1.1 Group Program . . . . . . . . . . . . . . . . . . . . . . . . 13 7.1.2 Group settings . . . . . . . . . . . . . . . . . . . . . . . . 14

7.2 Practical examples . . . . . . . . . . . . . . . . . . . . . . . . . . 15 7.2.1 The normal case - direct conversion between 2 formats . . 15 7.2.2 Prepare ASCII-text for word processing . . . . . . . . . . . 16 7.2.3 An example for a special problem . . . . . . . . . . . . . . 17

8 Known bugs 18 8.1 Internal bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 8.2 Problems with other programs . . . . . . . . . . . . . . . . . . . . 18

9 Future development 19

10 History 20

11 Credits 21

12 Address of the author 22

1 GENERAL

1 General

In spring 1996, when I released a little program called TextConv at the AmiNet, I did not expect to receive such a huge amount of reactions or to become published in an Amiga magazine (Germans Amiga Spezial 2/96). Because of these reactions I decided to continue the development of TextConv up to this new version.

Please excuse the long time that I needed for finishing this version. The causes are that I had to force my studies at the university because I didn‘t get any more money form the austrian gouvernment and the very long beta test time between the versions 1.50 and 1.51 because I tried to deliver a nearly bug-free program.

At this point, thanks to all the users of TextConv whose bug reports and suggestions helped me developing TextConv.

Please excuse the translation mistakes, notice that english is a foreign language for me.

2 INTRODUCTION

2 Introduction

Why an ASCII-text-converter?

• Did you ever try to edit a MS-DOS-text with a MS-Windows-editor?

• Or a Windows-text with OS/2?

• Or a Macintosh-text with an Amiga?

If you can answer one or more of these questions with YES, you already got the answer;-)

Well, the real cause is the following:

The ASCII-charset (ASCII=Amercian Standard Code For Information Interchange) isn‘t defined very exactly. Only the first 127 chars are equal for all computer systems, the rest varies from computer system to computer system (in this range are the most special chars and umlauts placed). Also the code for a line feed is different at the several system platforms:

Charsets:

ANSI ASCII: Amiga, UNIX, Windows
PC-850 ASCII: MS-DOS, OS/2^{1}
MAC ASCII: Apple Macintosh^{2}
ATARI ASCII: Atari^{3}

Linefeed code:

Amiga,UNIX:  LF
DOS,OS/2,Windows, Atari: CR+LF
Mac:  CR

When you want to transfer text files between these systems you have to bring the files into the right shape, and for this a converter like TextConv is needed.

It‘s important to know that this conversion usually causes losses; that means, as example, if a MS-DOS-text becomes converted to the Amiga format and than back to the MS-DOS-format, the new and the old file are not identically. The reason for this is that the Amiga does not have some specific PC-characters (graphics chars,for example). During conversion, this chars are converted to similar Amigachars, so that the original char is lost.

1The PC-850 charset is also known as IBM-charset

2The Macintosh-charset is similar to the ANSI-ASCII charset

3The Atari-charset is similar to the PC-850 charset

The same priciple is valid for Macintosh- or Atari-texts. If, for example, an Atari-Text should be converted to the Macintosh-format, it isn‘t useful to convert the text first into the Amiga-format (to read it with the Amiga or something else) and then convert it from the Amiga- to the Atari-format, because there are losses on the way from Atari to Amiga AND from Amiga to Macintosh. The conversion should always take the original file as source.

3 SYSTEM REQUIREMENTS, COMPATIBILITY

3 System requirements, compatibility

• 1 Amiga (or more :-) or the Amiga-emulator UAE

• Kickstart/Workbench 2.04 or higher

• Helvetica 11 - font in the FONTS:-directory

• about 230 kB free memory

• a minimum workbench screen hight of 256 pixels There shouldn‘t be any problemes if these requirements are fulfilled (please read the notices at the end of this section!). TextConv V1.51 has been tested with the following

3.1 Configurations

1) SCT Amiga2000C mode 1:

• Base Amiga 2000C, Kickstart/Workbench 2.04, 2.1, 3.1

• 2MB chip RAM, ECS Graphics via Multivision2000 flickerfixer

• Blizzard 2060/50, 28MB RAM, Blizzard SCSI-controller

• PicassoII gfxboard (original Picasso WB-emulation and Shareware-

CybergraphX-emulation)

2) SCT Amiga2000C mode 2:

• Base Amiga 2000C, Kickstart/Workbench 2.04, 2.1, 3.1

• 2MB chip RAM, ECS Graphics via Multivision2000 flickerfixer

• Original MC68000, 7.14MHz, BSC Tandem IDE HD

• FastRAM-configurations:

– 0MB FastRAM

– 2MB FastRAM from PicassoII with APM (Add Picasso Mem)

– 4MB FastRAM from GoldenGate486SLC/2 Bridgeboard with

     moremem

3) SCT UAE 0.68c for MS-DOS • Intel Pentium PC, 75/100/166MHz, 48MB RAM

• MS-DOS 7.0, Windows 95, OS/2 Warp4

• Kickstart/Workbench version 3.1 • UAE-configuration: 68000 and 68020-version, system requirements

from above must be fulfilled.

4) SCT UAE 0.68 Rel.14 for Windows

• Intel Pentium PC, 166MHz, 48MB RAM

• Windows95, DirectX 3.0

• Kickstart/Workbench version 3.1

• UAE-configuration: 68020-version, system requirements from above must be fulfilled.

5) MDS UAE 0.68c for MS-DOS

• Intel Pentium PC, 150MHz, 32MB RAM

• MS-DOS 7.0, Windows 95

• Kickstart/Workbench version 3.1

• UAE-configuration: 68000 und 68020-version, system requirements from above must be fulfilled.

3.2 Additional notices

Notice 1: The GUI is designed for WB-resolutions with a 4:3-proportion (e.g. Hires-Interlace, Productivity, (S)VGA-resolutions from GFX-boards). Such a mode isn‘t necessary, but it looks much better;-)

Notice 2: In case of an access to a CrossDos-drive the CrossDos-commodity (Note: the commodity, not the driver) may not be active, because this commodity also makes a (partial) conversion of textfiles like the way Text-Conv does. This leads to results that are not desired. Files that have been converted with the CrossDos-commodity usually can‘t become further processed via TextConv, because the conversion that CrossDos does is too inaccurate.

Notice 3: If text should be transfered to the Mac-emulator ShapeShifter via Clipboard, the format of these texts has to be Amiga-ACSII-format, because ShapeShifter converts text to the Mac-format itself when they are imported from the Amiga-clipboard.

4 WARRANTY, MAILWARE, COPYRIGHT, COPYING

4 Warranty, mailware, copyright, copying

4.1 Warranty

There is no warranty, neither for correct program functions nor for any damages caused by TextConv.

4.2 MailWare

Since version 1.50, TextConv is a MailWare-program. The principles for MailWare are the same as for FreeWare, that means, the usage of the program and the spreading is free. If you like the program and you use it regulary, you should send me an e-mail, a (picture) postcard or a letter. There you can tell me if you like the program or what you liked to have changed. So you can help me to find bugs and improve future versions of TextConv.

4.3 Copyright, copying

All rights are restricted to the author. \begin{matrix}c \\ ◯ \end{matrix}1995-97 SciTech SD. \begin{matrix}c \\ ◯ \end{matrix}1995-97 Alexander Haas

This program may be copied and spreaded as you like, but no fee may be charged in any form, except a suitable fee for copying, but max. the price of a DD-disk (20ATS/3DM/2US$) or, with printed manual, 50ATS/7DM/5US$.

TextConv consists of the following archives:

TextConv151d.lha: German version of TextConv V1.51 Contains the program, conversion tables, dockbrushes, Documentation in ASCII- and Amigaguide-format

TextConv151e.lha: English version of TextConv V1.51 Contains the program, conversion tables, dockbrushes, Documentation in ASCII- and Amigaguide-format

TextConvdocs151d.lha: German documentation in TeX-DVI and HTML-for-mat Contains the additional description of the ACT-conversiontable- fileformat in TeX-DVI-format.

TextConvdocs151e.lha: English documentation in TeX-DVI and HTML-for-mat Contains the additional description of the ACT-conversiontable- fileformat in TeX-DVI-format.

4 WARRANTY, MAILWARE, COPYRIGHT, COPYING

Each archive respectively its contents may only spread or published complete (compressed or uncompressed). This program may be used in any way (also commercial).

5 INSTALLATION

5 Installation

The installation is very easy, so I thought that it isn‘t necessary to write an Installerscript:

5.1 Standard installation process

Just copy the TextConv-folder into the desired target directory (at the Workbench via drag the folder, with CLI/Shell via the command COPY or with a DOStool like DOpus). No specific system setups (e.g. ASSIGN‘s) are necessary. But you have to be sure that the system requirements are fulfilled. Attention: the folder ConvTabs has to be placed into the same directory as the main program TextConv!

5.2 Additional installation notes

If you want to use TextConv via programs like ToolManager or Directory Opus please read the notes at the section Starting TextConv and the manuals of the respective program. Some dockbrushes e.g. for Toolmanager are included in this archive (directory Docks).

Toolmanager-settings:

ExecType: CLI
Stack:  4096
Priority:  0
Delay:  0
Arguments: activated

Be careful with the correct path names!! The setting Current Directory has to be set to the program folder, otherwise the conversion tables in the folder ConvTabs will not be found!!

DirectoryOpus-settings:

Type:  AmigaDOS

 Command line: [PFAD] TextConv {F}

({F} ...all selected files will be handed over)

Flags:  Run asynchronously

Workbench to front

(TextConv-window will always be opened at the WB-Screen)

6 STARTING OF TEXTCONV

6 Starting of TextConv

The usual starting of TextConv happens via Workbench or CLI:

Workbench: with double-click on the TextConv-icon

CLI/Shell: TextConv [[PATH] filename [PATH] filename ..]

There can be loaded max. 100 files at once.

Because of the CLI-interface the starting of TextConv via programs like Toolmanager or DirectoryOpus is recommended (see section Installation). When using Toolmanager, the desired files are just dragged to the respective dock; in DirectoryOpus, the desired files are marked first, then TextConv can be started by pressing the TextConv-button. Please notice that max. 100 files can be handed over!!

7 USING TEXTCONV

7 Using TextConv

After the start the TextConv-window will be opend. It‘s splitted into three groups:

1. Control buttons (group program)

2. Setting buttons (group settings)

3. Status line

If no files has been handed over at the start, the status line says

"No files loaded",

otherwise

"# files selected - START to convert".

IMPORTANT: Files that have been overhanded at startup or that have been ”loaded” at the ASL-request won‘t be really loaded, they just become marked for loading. The files will be loaded into memory during the conversion process. That‘s the cause why the settings can be choosen after the ”loading” process (but not after the start of the conversion process).

If you didn‘t overhand files at the startup, you have to press the Load files..- button. An ASL-requester appears, from where you can choose up to 100 files (multiselect by pressing the SHIFT-button; only possible, when all files are located in the same folder). This ASL-requester is also drag&drop-able, this means you can drag files form a Workbench-window itno the requester. This causes that you change the actual directory to the one you dropped into the requester. Please notice that only the first marked file will be overhanded!

When using the program ArcHandler its also possible to convert files directly out of LHA-archives. At the moment only reading from archives is possible, writing NOT!

After the choosing of the desired files you have to tell the program the format of the source files and into which format they should be converted. This happens with the following cycle-gadgets (see also Description of the gadgets):

• Read files as.. (format source file(s))

• Save files as.. (format target file(s))

• Conversion mode (with or without contol chars respectively remove control chars or convert only control chars; usually always Standardconversion, this means convert chars and control chars; see also Practical examples)

• Save mode (how to save the files)

Press the start button to start the conversion process after you have done your desired settings. Depeding on the kind of Save mode you might be asked about the target directory or -file names.

If a file exists with the same name as the one that TextConv wants to write, a requester appers and asks whether the file should become overwritten or a new name should be set. Notice that the file that actually becomes converted can‘t be overwritten (this means replacing of a file is not possible). The gadgets and menus are switched off during the conversion process to avoid unwanted inputs.

The status of the conversion process is displayed at the status line, the name of the actual file and the conversion in percent is visible. If everything has been alright, the message File(s) converted is displayed at the status line, otherwise an error message will be shown.

After a short period all files have been converted and now you can convert the same files with other settings, you can load new files or you can exit the program by pressing the Quit button. After pressing the Quit-button you will be asked whether you really want to finish or not.

7.1 Description of the gadgets

7.1.1 Group Program

Load files: Used for choosing the files to load; directory can be choosen by drag&drop, multiselect by pressing the SHIFT-button; at multiselect all files have to be located at the same directory (this is a restriction caused by the ASL-requester, when overhanding files via CLI-interface this restriction doesn‘t exist). A maximum of 100 files (any size) can be overhanded.

Start: Starts the conversion process. During the conversion all gadgets are switched off.

About TextConv: Opens a window containing general informations about Text-Conv.

Quit: Quits TextConv. Before quitting a requester appears that asks whether finishing really or not.

All the points above are also accessable via the Project-menu.

7.1.2 Group settings

Read files as: Sets the source format.

Possible settings are:

• Amiga ASCII (UNIX)

• MS-DOS/OS2-ASCII (IBM/CP850)

• MS-Windows ANSI

• Macintosh

• Atari

Save files as: Sets the target format.

Possible settings are:

• Amiga ASCII (UNIX)

• MS-DOS/OS2-ASCII (IBM/CP850)

• MS-Windows ANSI

• Macintosh

• Atari

Conversion mode: Sets the conversion mode for the control chars. Possible settings are:

• Standardconversion Control chars and normal chars will be converted. Use: this is the standard setting

• Don‘t convert control chars Normal chars will be converted, control chars will be not changed. Use: for special problems

• Remove control chars Normal chars will be converted, control chars will be removed. Use: preparing the text for later use with a word processor (for automatic justified lines or something similar) ATTENTION: indentions will also be removed!

• Convert only control chars Normal chars will not be changed, control chars will be converted. Use: for special problems

Save mode: Sets the way that converted files should be saved.

Possible settings are:

• each file seperate Useful for a small number of files For each file the target directory and the target name is asked (the original name is NOT overhanded automatically!).

• same dir. another extension In this case the source- and target directory are the same; the converted files are marked with a special extension that depends on the target format: Possible extensions are:

– .ami...Amiga or Unix-format

– .dos...MS-DOS or OS/2-format

– .win...MS-Windows-format

– .mac...Apple-Macintosh-format

– .ata...Atari-format

• into another directory All files will be written into a specified target dirctory; the filenames won‘t be changed.

7.2 Practical examples

7.2.1 The normal case - direct conversion between 2 formats

Problem description: A MS-DOS ASCII-text should be converted to the Amigaformat.

Solution: The settings for this case are:

• Read files as: IBM/MSDOS/OS2 ASCII

• Save files as: Amiga ASCII (Unix)

• Conversion mode: Standardconversion

• Save mode: (any)

The text can now be used with any Amiga text editor.

Problem description: Some Atari-texts should be converted to Mac-format; the texts should be viewed at the Amiga before processed at the Mac.

Solution: Because of the problem that the conversion causes losses, the way Atari → Amiga → Mac should be avoided. It‘s better to create both the Amiga and the Macintosh files from the original sourcefile. After the file selection, we first create the Amiga files:

• Read files as: Atari ST

• Save files as: Amiga ASCII (Unix)

• Conversion mode: Standardconversion

• Save mode: same dir. another extension

Press the Start-button to start the conversion. The source folder now contains also the converted files, marked with the extension .ami. Because TextConv remembers the source files also after the conversion, you just have to change the setting

• Save files as: Apple Macintosh;

you don‘t have to select the source files again. After pressing the Startbutton the Macintosh-files will be created (Extension .mac). If you use the setting same dir. another extension, be careful that the max. allowed number of chars for filenames (Amiga: 25 or 30) doesn‘t exceed.

7.2.2 Prepare ASCII-text for word processing

Problem description: An Amiga-ASCII-text should be handed over to the Amiga word-processor FinalCopy, so that flushed line justification is possible.

Solution: To create flushed text, the lines may not be ended with line feed. When a word processor detects such a line feed, it‘ s interpreted as an indent. To avoid this, the LF-codes have to be removed. Because FinalCopy has no import filter that removes line feeds automatically, a converter like TextConv has to do this job. The settings for this problem are:

• Read files as: Amiga ASCII (Unix)

• Save files as: Amiga ASCII (Unix)

• Conversion mode: Remove control chars

• Save mode: (any)

Notice that indents and empty lines also will be removed (this may be changed at the next release).

Problem description: A Mac-ASCII-text should be handed over to MS-Win-word, so that flushed line justification is possible.

Solution: In contrast to FinalCopy Winword has a text import filter that is able to remove linefeeds (but difficulty to use depends on the respective WinWord-version). Because of that, that a conversion has to be done via TextConv from Mac to PC-ASCII anyway, the removing of the LF‘s can be done in the same step. The settings are:

• Read files as: Apple Macintosh

• Save files as: MS Windows ANSI

• Conversion mode: Remove control chars

• Save mode: (any)

The file can now be imported into WinWord via the Text only import filter. Attention: also indets will be removed.

7.2.3 An example for a special problem

Problem description: An e-mail has been received on an OS/2-PC, the contents of the mail is a special Amiga-ASCII-text (concrete: a registration file), which has to be saved from an OS/2-editor (via copy&paste). This transfer causes a changing of the control chars form LF to CRLF.

Solution: The normal chars are neither changed from the transmission via clipboard nor from the saving process of the editor; only the control chars will be changed from LF to CRLF. The text is still in Amiga-ASCII-code; the control chars are in MS-DOS-format. The correct setting to solve this problem are:

• Read files as: IBM/MSDOS/OS2 ASCII

• Save files as: Amiga ASCII (Unix)

• Conversion mode: Convert only control chars

• Save mode: (any)

With the setting Convert only control chars only the changing from LF to CRLF is undone, all normal chars are not changed.

8 KNOWN BUGS

8 Known bugs

Not one program all over world is 100% bug-free, TextConv is no exception. Following bugs and problems are known and will be removed (I hope so) at the next release:

8.1 Internal bugs

• In very rare cases TextConv thinks that a file already exists (Warning before overwriting) although this isn‘t TRUE.

8.2 Problems with other programs

• When overhanding files from ToolManager it may happen, that TextConv starts with the status message "No files loaded". This should not be caused by TextConv, but from a restriction of the maximum length of a CLI parameter line from ToolManager.

• When using TextConv with UAE 0.68c for MS-DOS, there can be problems with direct (write) access to virtual DOS-disks; especially when UAE has been started from Win95. This should be a bug in UAE, because other programs have the same problem.

9 FUTURE DEVELOPMENT

9 Future development

The basic development of TextConv is finished with the release of this version. The following changes are planned:

• Possibility to keep indents at the setting Remove control chars

• Changing of the file management, so that the last choosen directory appears automatically at the next load process.

• Bug-cleaning (if there are still any;-)

• Maybe supporting of 1 or 2 more charsets

• Additional conversion parameters A planned adaption to a MUI-interface will not be developed. The cause is that many users use TextConv with UAE, and MUI is very slow under this conditions. \underline{Eventually} planned is a Windows-version of TextConv.

10 HISTORY

10 History

V1.51d/e: Second official AmiNet-release • CyberGuard- and EnforcerHit-free

• Bug removed at the ASL-requester-management that sometimes caused a system crash.

• english version V1.50d: Inofficial release, available on request

• nearly complete new code

• partial modular concept; conversion tables as external files

• Use of a read-/write-buffer; this causes 1. Size of the convertable files only limited by hard disk space

2. About 5-20 times faster

• Implementation of the Atari-format

• Very ugly bug at the Macintosh-conversion removed (uncorrect char conversion at the ASCII-range 127-255)

• New conversion modes:

1. Remove control chars

2. Convert only control chars

• Changing the program classification from FreeWare to MailWare V1.2-1-4: Internal version numbers V1.10d/e: Inofficial release, available on request

• Warning about overwriting of files

• Dynamic memory management, this causes that the max. size of the convertable files is only limited by the amount of free memory

• Small change at the GUI, so that the whole window is visible at screens with a hight of 256 pixels.

• Removing of some little bugs (correct on/off-switching of the gadgets)

• english version V1.00d: First official AmiNet-release

Awarded from the german Amiga-magazine AmigaSpezial 2/96

11 CREDITS

11 Credits

Used Copyrights:

DirectoryOpus: \begin{matrix}c \\ ◯ \end{matrix}Jonathan Potter, GPSoftware
ToolManager:  \begin{matrix}c \\ ◯ \end{matrix} Stefan Becker
MUI:  \begin{matrix}c \\ ◯ \end{matrix} Stefan Stunz, SASG
ArcHandler:  \begin{matrix}c \\ ◯ \end{matrix} Rafael D‘Halleweyn
AddPicassoMem: \begin{matrix}c \\ ◯ \end{matrix} Nicola Soggia
UAE:  \begin{matrix}c \\ ◯ \end{matrix} Bernd Schmidt
UAE Dos:  \begin{matrix}c \\ ◯ \end{matrix} Gustavo Goedert
UAE Windows: \begin{matrix}c \\ ◯ \end{matrix} Mathias Ortmann
GoldenGate:  \begin{matrix}c \\ ◯ \end{matrix} Vortex
Amiga, Workbench, CLI, Shell: \begin{matrix}c \\ ◯ \end{matrix} Amiga International

The programs ToolManager, MUI, ArcHandler und AddPicassoMem can be downloaded from the AmiNet.

12 ADDRESS OF THE AUTHOR

12 Address of the author

Bug reports, suggestions, donations, letters, post cards etc to:

SciTech Software Development Haas Alexander Mitterndorf 150 A-4801 Traunkirchen AUSTRIA

e-mail: scitech@ping.at
WWW: http://members.ping.at/scitech